# Anosov diffeomorphisms of open surfaces

### v1

@@ 4.4. Global Holonomy Invariance of the Margulis Measures
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}(\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|\pm 2)}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
Note however, that for $k$ sufficiently large that $f^{-n}(x_{k},\omega_{k})$ intersects neither none of the sets $\mathcal{D}_{N}^{R},\mathcal{D}_{N}^{S}$ or $\mathcal{D}_{N}^{T}$. Thus we see in fact this stronger estimate:
(3) $\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}(\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|)}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
$\mu_{\underline{R}}(\{x\})+\mu_{\underline{T}}(\{x\})=\inf_{N}\sum_{[C]\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}([C])+\sum_{[C]\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}([C]).$
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\mu_{\underline{R}}(\{x\})+\mu{\underline{T}}(\{x\}).$
@@
The proof for the remaining cases is identical, but the notation is more or less complicated depending on the particular configuration of the Markov rectangles. One takes a cover of the set being measured by the images of cylinders and then estimates the measures of the cylinders by looking at intersections exactly as before.
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\leq n_{k}}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(\underline{a},a_{0}))\right|}{\sum_{i\leq n_{k}}e^{-ih_{G}(\sigma)}\left|W^{u}(\underline{a})\cap f^{-i}(W^{s}(\underline{a},a_{0}))\right|}.$
Recall that in Corollary 4.16, we defined measures on each global leaf. However, for the leaves that the above Theorem is silent on, we will now redefine the Margulis measures on them by using local holonomies. The above theorem asserts that this can be done in a consistent way.
@@ 5. Application to Anosov diffeomorphisms on Open Surfaces
In this section, we give an application of the construction of the Margulis measures to Anosov diffeomorphisms on open surfaces under some assumptions. In the first subsection, we show that the Margulis measures are compatible with global product structure and give a simple description of the universal cover of $M$ when $f$ has product structure. In the second subsection, we show that this product structure forces $M$ to be $\mathbb{T}^{2}$.
5.1. Margulis measures and product structure
In this section we show that the Margulis measures give a particularly nice description of the universal cover of $M$ when it has product structure.
First we recall the definition of product structure for an Anosov diffeomorphism $f\colon M\to M$. If we let $\widetilde{W}^{u}$ and $\widetilde{W}^{s}$ denote the stable and unstable foliations of the universal cover of $M$, then we say that $f$ has global product structure, or briefly, product structure, if for any $x,y\in M$, $\left|\widetilde{W}^{u}(x)\cap\widetilde{W}^{s}(y)\right|=1$, i.e. every pair of stable and unstable leaves meets in exactly one point. In the case of a surface $M$, this implies that the foliations on $\widetilde{M}\approx\mathbb{R}^{2}$ look like the foliations of $\mathbb{R}^{2}$ by vertical and horizontal lines.
@@ 5.1. Margulis measures and product structure
Suppose that $f\colon M\to M$ is a uniform Anosov diffeomorphism with product structure. Let $\mu^{u}$ be a system of Margulis measures from Corollary 4.33. For a fixed point $p$ of $f$, we will define a map $\pi_{p}$ by
@@
Let $\widetilde{M}$ denote the universal cover of $M$ and $\widetilde{W}^{s}$ and $\widetilde{W}^{u}$ be the lifts of the stable and unstable foliations of $f$ to $\widetilde{M}$. Let $P\colon\mathbb{R}\times\mathbb{R}\to\widetilde{M}$ be the map that shows that $\widetilde{W}^{u}$ and $\widetilde{W}^{s}$ have product structure in the universal cover with the point $(0,0)$ mapping to some point in $\pi^{-1}(p)$. Then clearly $\pi\circ P$ is the universal cover of $M$.
It then suffices to check that this map $P$ satisfies the property in (5). Certainly, due to the full support of $\mu^{u}$ and $\mu^{s}$, we that as long as it exists any map satisfying (5) is uniquely defined up to the choice of basepoint $p$. Note now that the property (5) is satisfied within any foliation box due to the holonomy invariance of the Margulis measures. The global product structure shows that for any $x,y,z$ as above, we may include them into the image of a single large foliation box in the universal cover, and the result follows. ∎
Remark 5.2.
The proof of Proposition 5.1 makes use of the product structure. A posteriori, it follows that for every leaf of $W^{u}(x)$ that $W^{u}(x)$ has infinite measure with respect to $\mu^{u}$. In fact, if we had known that every leaf had infinite volume a priori, then we could have deduced that there was global product structure for these foliations along similar lines. This is the argument of Hiraide [Hir01] that codimension 1 Anosov diffeomorphisms have global product structure. The application in this paper could also be concluded if we had assumed instead that every leaf had infinite measure. Note that if we tried to define the universal cover by (5) but every leaf did not have infinite measure, then we would not necessarily been able to define the map: consider for instance the holonomy between an infinite measure leaf and a finite measure leaf.
5.2. Application to uniform Anosov diffeomorphisms
Using the universal cover constructed in the previous section we obtain an application showing that a type of uniform Anosov diffeomorphism on an open surface cannot exist.
Theorem 5.3.
Suppose that $f\colon M^{2}\to M^{2}$ is a uniform Anosov diffeomorphism of a surface such that periodic points are dense, and the unstable and stable manifolds of each periodic point are dense, and having product structure. Then $M^{2}$ is closed.
@@ 5.2. Application to uniform Anosov diffeomorphisms
For the sake of contradiction, suppose that $M$ is open. The proof has two parts. In the first part, we use the Margulis measures and the argument of Hiraide [Hir01] to show that $\pi_{1}(M)$ is abelian, which implies that either $\pi_{1}(M)$ is trivial or $\mathbb{Z}$. In the second part, we show that such a surface cannot admit a uniform Anosov diffeomorphism by considering its dynamically invariant foliations.
Part 1: $\pi_{1}(M)$ is abelian. From the previous subsection, we have that the map
$\pi_{p}\colon W^{u}(p)\times W^{s}(p)\to M$
is the universal cover of $M$.
We now lift $f$ to the universal cover by $\pi$ so that $(p,p)$ is fixed. Define the lift by $F=f|_{W^{u}(p)}\times f|_{W^{s}(p)}$. It follows that this is a lift of $f$ by the invariance of the stable and unstable foliations: note that $F$ is certainly a lift of $f$ when restricted to $W^{u}(p)\times\{p\}$ and $\{p\}\times W^{s}(p)$; hence as $F$ preserves the product structure on $\widetilde{M}$ and the stable and unstable foliations on $M$, $F$ is necessarily the lift of $F$ as it is a lift on these two distinguished leaves, $W^{u}(p)\times\{p\}$ and $\{p\}\times W^{s}(p)$.
We now study the fundamental group by viewing it as the group of all deck transformations of $\pi_{p}\colon\widetilde{M}\to M$. Let $\alpha\in\pi_{1}(M)$ and $x\in W^{u}(p)\times W^{s}(p)$. Let $\widetilde{\alpha}(x)\in W^{u}(p)$ denote the projection of $\alpha(x)$ onto $W^{u}(p)$ along the $W^{s}(p)$ fibers. It follows that $\mu^{u}(I)=\mu^{u}(\widetilde{\alpha}(I))$ because if $\pi_{p}(I)=\pi_{p}(J)$, then $\mu^{u}(\pi_{p}(I))=\mu^{u}(\pi_{p}(J))$. Hence as $\pi_{p}(I)=\pi_{p}(\alpha(I))$, it follows that $I$ and $\alpha(I)$ have the same $\mu^{u}$ measure. Thus we see that $\widetilde{\alpha}\colon W^{u}(p)\to W^{u}(p)$ may be regarded as a measure preserving transformation of $W^{u}(p)$ endowed with $\mu^{u}_{W^{u}(p)}$.
We claim that the map $\alpha\mapsto\widetilde{\alpha}$ is an injective homomorphism. Note that $\alpha$ preserves the product structure on $\widetilde{M}$ as it is the lift of the stable and unstable foliations on $M$. Hence projection onto the fixed $W^{u}(p)$ leaf commutes with applying deck transformations, and hence the map $\widetilde{\alpha}\mapsto\alpha$ is a homomorphism. For injectivity, suppose that $\widetilde{\alpha}$ is the identity and suppose $q\in W^{u}(p)$. This implies that $\{\widetilde{\alpha}\colon\alpha\in\pi_{1}(M)\}$ is a free abelian group. Note that if the restriction $\widetilde{\alpha}\colon W^{u}(p)\to W^{u}(p)$ is trivial, then $\alpha$ carries $W^{s}(p)$ to itself. Hence $\pi_{p}\circ\alpha=\pi_{p}$ when restricted to $W^{s}(p)$. But as $W^{s}(p)$ is simply connected and $\pi_{p}|_{W^{s}(p)}$ is a cover, we see that $\alpha$ is trivial restricted to $W^{s}(p)$. Thus we see that $\widetilde{\alpha}\mapsto\alpha$ is an injection.
We have just embedded $\pi_{1}(M)$ as a subgroup of $\text{Isom}(\mathbb{R})$ as for any other measure $(\mathbb{R},\mu)$ if $\mu$ is fully supported its group of measure preserve transformations embeds in $\text{Isom}(\mathbb{R})$. The fundamental group of an open surface is free [Sti93, Sec. 4.2.2]. If $\pi_{1}(M)$ contains an element that is not a translation, then this element is torsion, which is impossible in a free group. Hence $\pi_{1}(M)$ actually embeds as a group of translations of $\mathbb{R}$, hence it abelian. Thus as $\pi_{1}(M)$ is abelian it is trivial or $\mathbb{Z}$.
@@
Part 2: $\pi_{1}(M)$ is not trivial or $\mathbb{Z}$. We give a different argument in each case. In both cases we find that the manifold has insufficient topology to allow $f$ to be Anosov. In the case that $\pi_{1}(M)$ is trivial, $M$ is diffeomorphic to $\mathbb{R}^{2}$. Hence the result of Mendes [Men77, Prop. 1.2] implies that $f$ has at most one non-wandering point; but this is impossible because the periodic points of $f$ are dense. If $\pi_{1}(M)=\mathbb{Z}$, then $M$ is diffeomorphic to a cylinder or a Möbius strip. However, [HHU19, Lemma 4.3] shows that the cylinder does not support Anosov dynamics with dense periodic points. A similar argument to the one in [HHU19, Lemma 4.3] shows that Anosov dynamics with dense periodic points and dense $\mathcal{W}^{u/s}$ foliations are not possible on the Möbius strip. Thus we have exhausted all possibilities for the topology of $M$ and are done. ∎

### v2

@@ 3.5. Global Holonomy Invariance of the Margulis Measures
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}(\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|\pm 2)}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}_{0}(\underline{a}))\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
Note, however, that for $k$ sufficiently large that $f^{-n}(x_{k},\omega_{k})$ intersects none of the sets $\mathcal{D}_{N}^{R},\mathcal{D}_{N}^{S}$ or $\mathcal{D}_{N}^{T}$. Thus we see in fact this stronger estimate:
(4) $\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}(\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|)}{\sum_{i\geq 0}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}_{0}(\underline{a}))\cap f^{-i}(W^{s}(x_{k},\omega_{k}))\right|}.$
@@
$\mu_{\underline{R}}(\{x\})+\mu_{\underline{T}}(\{x\})=\inf_{N}\sum_{C\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}(C)+\sum_{C\in\mathcal{C}_{N}^{R}(\{x\})}\widehat{\mu}_{\underline{R}}(C).$
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\mu_{\underline{R}}(\{x\})+\mu_{\underline{T}}(\{x\}).$
@@
The proof for the remaining cases of the relative positions of the Markov rectangles is identical, but the notation is more or less complicated depending on the particular configuration of the rectangles. One takes a cover of the set being measured by the images of cylinders and then estimates the measures of the cylinders by looking at intersections exactly as before.
@@
$\mu_{\underline{S}}(\{\Gamma(x)\})\leq\lim_{k\to\infty}\frac{\sum_{i\leq n_{k}}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(\mathcal{D}_{N}^{R}\cup\mathcal{D}_{N}^{T})\cap f^{-i}(W^{s}(\underline{a},a_{0}))\right|}{\sum_{i\leq n_{k}}e^{-ih_{G}(\sigma)}\left|\widehat{\pi}(W^{u}(\underline{a}))\cap f^{-i}(W^{s}(\underline{a},a_{0}))\right|}.$
Recall that in Corollary 3.15, we defined measures on each global leaf. However, for the non-proper leaves we will now redefine the Margulis measures by using local holonomies. The above theorem asserts that this can be done in a consistent way due to the continuity and consistency of the holonomies.
@@ 4. Application to Anosov Diffeomorphisms on Open Surfaces
In this section, we give an application of the construction of the Margulis measures to Anosov diffeomorphisms on open surfaces under some assumptions. In the first subsection, we show that the Margulis measures are compatible with global product structure and give a simple description of the universal cover of $M$ when $f$ has global product structure.
4.1. Margulis Measures and Global Product Structure
In this section, we show that the Margulis measures give a particularly nice description of the universal cover of $M$ when it has Global product structure.
First we recall the definition of global product structure for an Anosov diffeomorphism $f\colon M\to M$. If we let $\widetilde{W}^{u}$ and $\widetilde{W}^{s}$ denote the stable and unstable foliations of the universal cover of $M$, then we say that $f$ has global product structure if for any $x,y\in M$, $\left|\widetilde{W}^{u}(x)\cap\widetilde{W}^{s}(y)\right|=1$, i.e. every pair of stable and unstable leaves meets in exactly one point. In the case of a surface $M$, this implies that the foliations on $\widetilde{M}\approx\mathbb{R}^{2}$ look like the foliations of $\mathbb{R}^{2}$ by vertical and horizontal lines.
@@ 4.1. Margulis Measures and Global Product Structure
Suppose that $f\colon M\to M$ is a uniform Anosov diffeomorphism with global product structure. Let $\mu^{u}$ be a system of Margulis measures from Corollary 3.32. For a fixed point $p$ of $f$, we will define a map $\pi_{p}$ by
@@
Let $\widetilde{M}$ denote the universal cover of $M$ and $\widetilde{W}^{s}$ and $\widetilde{W}^{u}$ be the lifts of the stable and unstable foliations of $f$ to $\widetilde{M}$. Let $P\colon\mathbb{R}\times\mathbb{R}\to\widetilde{M}$ be the map that shows that $\widetilde{W}^{u}$ and $\widetilde{W}^{s}$ have global product structure in the universal cover with the point $(0,0)$ mapping to some point in $\pi^{-1}(p)$. Then clearly $\pi\circ P$ is the universal cover of $M$.
It then suffices to check that this map $P$ satisfies the property in (6). Certainly, due to the full support of $\mu^{u}$ and $\mu^{s}$, as long as it is defined any map $\pi_{p}$ satisfying (6) is uniquely defined up to the choice of basepoint $p$. Note now that the property (6) is satisfied within any foliation box due to the holonomy invariance of the Margulis measures. The global product structure shows that for any $x,y,z$ as above, we may include them into the image of a single large foliation box in the universal cover, and the result follows. ∎
The proof of Proposition 4.1 makes use of the global product structure. A posteriori, it follows that for every leaf of $W^{u}$ that $W^{u}(x)$ has infinite measure with respect to $\mu^{u}$. In fact, if we had known that every leaf had infinite volume a priori, then we could have deduced that there was global product structure for these foliations along similar lines. This is the argument of Hiraide [Hir01] that codimension 1 Anosov diffeomorphisms have global product structure. The application in this paper could also be concluded if we had assumed instead that every leaf had infinite measure. Note that if we tried to define the universal cover by (6) but every leaf did not have infinite measure, then we would not necessarily have been able to define the map: consider for instance the holonomy between an infinite measure leaf and a finite measure leaf.
We say that a foliation $\mathcal{F}$ of a manifold $M$ is $\mathbb{R}$-covered when the lift of $\mathcal{F}$ to the universal cover $\widetilde{M}$ of $M$ has leaf space isomorphic to $\mathbb{R}$. See, e.g. [Fen94, p. 82].
Proposition 4.2.
Suppose that $f\colon M\to M$ is a uniform Anosov diffeomorphism and that $f$ has Margulis measures $\mu^{u}$ as in Theorem 1.5 that assign infinite measure to each unstable leaf. Then $W^{s}$ is $\mathbb{R}$-covered.
@@
The proof follows the ideas of Hiraide. There are some minor complications only because the measures we consider might have atoms. We can construct a covering map analogously with the construction in Proposition 4.1.
Take a fixed point $p$ of $f$ and leaves $W^{s}(p)$ and $W^{u}(p)$. Then we define a map $\pi\colon W^{s}(p)\times W^{u}(p)\to M$ by $\pi(z,w)=q$ exactly when $q$ is the point in $W^{u}(z)$ such that $\mu^{u}([p,w])=\mu^{u}([z,q])$ and this agrees with the orientation where $W^{u}(z)$ meets $W^{s}(p)$ at $z$. We must now check that there is such a point $q$. Note next that for any $z\in W^{s}(p)$ there is a product neighborhood containing $z$ whose $W^{u}$ plaques are arbitrarily long. Hence locally the definition of $\pi$ makes sense; what we need to do is check that we can extend this definition from an initial product neighborhood containing $p\in W^{s}(p)$. Fix some $r>0$ and $M>0$. Then, using that the Margulis measure of each $W^{u}(p)$ leaf is infinite, for any $z\in B_{r}(p)\subseteq W^{s}(p)$, we may find a product neighborhood of $z$ containing all points $q^{\prime}$ with $\mu^{u}([w,q^{\prime}])\leq M$ where $w$ is the corresponding point in $W^{s}(p)$. These neighborhoods give an open cover of $B_{r}(p)$. By compactness of $B_{r}(p)\subseteq W^{s}(p)$, we can pass to a finite subcover of this set. Holonomy invariance of the Margulis measure then gives the needed conclusion on this subset. Thus the map $\pi$ is well defined.
We now check that this map defines the universal cover. We must check that $\pi$ is a continuous, surjective, local homeomorphism. Essentially all of these points follow from working in a small neighborhood of a product neighborhood $(z,w)\in W^{s}(p)\times W^{u}(p)$. Continuity is immediate from the full support of the Margulis measures. The map is a local injection because it is increasing with respect to the Margulis measures along $W^{u}$ leaves. Hence due to invariance of domain, it is a local homeomorphism as it is a continuous, local injection. Finally, the map is a surjection because $W^{u}(p)$ is dense by assumption as $p$ is a periodic point; surjectivity then follows from the uniform transversality of the stable and unstable foliations. ∎
Once the stable foliation is known to be $\mathbb{R}$-covered, one can conclude that the fundamental group of $M$ is abelian. Thomas Barthelme pointed out the following proof to us, which he told us was also known to Jana Rodriguez Hertz and Raul Ures.
Proposition 4.3.
Suppose that $M$ is a closed or open surface admitting a foliation $\mathcal{F}$ that is $\mathbb{R}$-covered and has no closed leaves, such as $\mathcal{W}^{s}$. Then $M$ has abelian fundamental group.
Proof.
Consider the lift of $\mathcal{F}$ to $\widetilde{M}$, $\widetilde{\mathcal{F}}$. Now consider the action of $\pi_{1}(M)$ on the leaf space of $\widetilde{\mathcal{F}}$. Note that the action must be free: otherwise $\mathcal{F}$ would contain a closed leaf. As $\pi_{1}(M)$ acts freely on $\mathbb{R}$, it follows that $\pi_{1}(M)$ is an Archimedean group. But by Hölder’s theorem [Ghy01, Thm. 6.10], Archimedean groups are abelian. ∎
In particular Theorem 1.6 now follows.
Proof of Theorem 1.6..
@@
From Proposition 4.2 and Proposition 4.3, it follows that $\pi_{1}(M)$ is abelian. From the classification of open surfaces [Sti93, § 4.2.2], this implies that $M$ is either $\mathbb{R}^{2}$ or $\mathbb{S}^{1}\times\mathbb{R}$, or a Möbius strip. In the case of $\mathbb{R}^{2}$, the result of Mendes [Men77, Proposition 1.2] implies that $f$ has at most one non-wandering point; but this is impossible because the periodic points of $f$ are dense. If $M$ is diffeomorphic to a cylinder [HHU19, Lemma 4.3] shows that the cylinder does not support Anosov dynamics with dense periodic points. A similar argument to the one in [HHU19, Lemma 4.3] shows that Anosov dynamics with dense periodic points and dense $\mathcal{W}^{u/s}$ foliations are not possible on the Möbius strip. Thus we have exhausted all possibilities for the topology of $M$ and are done. ∎